# A Simple Algorithm for Trimmed Multipoint Evaluation

### v1

@@ 3.2 Trimmed Multipoint Evaluation
For the proof of Theorem 3.1, consider the algorithm TrimmedEvaluation. Throughout we assume that the algorithm has access to the grid $Z$ and for simplicity, we omit $Z$ in the recursive calls.
@@
For $n=0$, the polynomial $P$ does not depend on any variable. Thus, we correctly return the constant $P$. For $n\geq 1$, we write $P(X_{1},\dots,X_{n})=\sum_{i=0}^{d}P_{i}(X_{1},\dots,X_{n-1})\cdot X_{n}^{i}$ to obtain $d+1$ many $(n-1)$-variate polynomials $P_{0},\dots,P_{d}$ of degrees $D,\dots,D-d$, respectively. Formally, each polynomial $P_{i}$ is the coefficient of $P$ viewed as a univariate polynomial in $X_{n}$. Let $V\coloneqq V(z_{n,0},\dots,z_{n,d})$ be a Vandermonde matrix. It is well-known that $V$ is invertible if and only if all grid points $z_{n,i}$ are distinct, and in this case the LU decomposition $V=L\cdot U$ in Line 1 always exists (see e.g. [19]). Further, let $Q_{0},\dotsc,Q_{d}$ as computed in Line 8 of TrimmedEvaluation. Note that each polynomial $Q_{j}(X_{1},\dotsc,X_{n-1})=\sum_{i=j}^{d}U_{j,i}\cdot P_{i}(X_{1},\dotsc,X_{n-1})$ has degree $D-j$. Indeed, the upper triangular structure of $U$ guarantees that $Q_{j}$ has at most degree $D-j$, because the first non-zero entry in the $j$-th row of $U$ corresponds to the polynomial $P_{j}$ of degree $D-j$ while all other $P_{i}$’s with (possible) non-zero coefficients have a smaller degree.
Calling TrimmedEvaluation$(Q_{j})$, we recursively compute the evaluations $Q_{j}(Z_{\ell^{\prime}})$ for all $\ell^{\prime}\in\binom{[n-1]}{\leq D-j}_{d}$ in Lines 1 and 1. Next, we focus on any iteration $\ell^{\prime}$ of the loop in Line 1. Let $k=\min\{d,D-\sum_{j=1}^{n-1}\ell^{\prime}_{j}\}$, then we show that we correctly compute the evaluations $P(Z_{\ell})$. For each $j\in\{0,\dotsc,k\}$, we have
@@
$=P\left\lparen Z_{(\ell^{\prime},j)}\right\rparen,$
where the first equality follows from the fact that $L_{j,i}=0$ whenever $i>k\geq j$. Notably, restricting the computation to $k$ is crucial to ensure that we only use the evaluations $Q_{j}(Z_{\ell^{\prime}})$ that we actually computed. This proves that TrimmedEvaluation correctly computes the evaluations $P\left\lparen Z_{(\ell^{\prime},0)}\right\rparen,\dotsc,P\left\lparen Z_{(\ell^{\prime},k)}\right\rparen$ ) in Line 13. Since each $Z_{\ell}$ can be expressed as some $Z_{(\ell^{\prime},j)}$ as above, the loop in Line 1 recovers all evaluations $P(Z_{\ell})$. As a result, the algorithm TrimmedEvaluation correctly computes all evaluations $P(Z_{\ell})$.
@@
Let $T(n,d,D)$ be the running time of TrimmedEvaluation. The time to compute the LU decomposition in Line $4$ and the matrix-vector products in Lines $5$ and $9$ can be bounded by $\operatorname{poly}(n,d)$. Consequently, the running time of the algorithm without the recursive calls in Line $7$ is $\binom{n}{\leq D}_{d}\cdot M(n,d)$ for some function [^6] In our context, we are content with bounding $M(n,d)=\operatorname{poly}(n,d)$; however, let us remark that following van der Hoeven and Schost [25] we could achieve a dependence on $d$ that is only polylogarithmic. The main insight is that the L- and U-factors of a Vandermonde matrix are sufficiently structured to support matrix-vector operations in time $O(d(\log d)^{O(1)})$. $M(n,d)=\operatorname{poly}(n,d)$. Thus, the algorithm admits the following recurrence
@@
Combining Lemma 3.2 and Lemma 3.4 concludes the proof of Theorem 3.1.
@@
5 Compute the LU decomposition $V(z_{n,0},\dots,z_{n,d})^{-1}=L\cdot U$.
6 for $\ell^{\prime}\in\binom{[n-1]}{\leq D-j}_{d}$ do
@@ 7       Let $k=\min\{d,D-\sum_{j=1}^{n-1}\ell^{\prime}_{j}\}$ and compute all evaluations
$\hskip 35.56593pt\begin{pmatrix}\beta_{(\ell^{\prime},0)}\\ \vdots\\ \beta_{(\ell^{\prime},k)}\\ \end{pmatrix}=\begin{pmatrix}U_{0,0}&\cdots&U_{0,k}\\ &\ddots&\vdots\\ &&U_{k,k}\\ \end{pmatrix}\cdot\begin{pmatrix}\alpha_{(\ell^{\prime},0)}\\ \vdots\\ \alpha_{(\ell^{\prime},k)}\\ \end{pmatrix}.\;$
@@
10       Recursively call TrimmedInterpolation($\left\lparen(\beta_{(\ell^{\prime},j)})_{\ell^{\prime}}\right\rparen$ ) for ${\ell^{\prime}}\in\binom{[n-1]}{\leq D-j}_{d}$ to interpolate the $(n-1)$-variate polynomial $Q_{j}$.
@@
$\hskip 65.44142pt\begin{pmatrix}P_{0}\\ \vdots\\ P_{d}\\ \end{pmatrix}=\begin{pmatrix}L_{0,0}&&\\ \vdots&\ddots&\\ L_{d,0}&\cdots&L_{d,d}\\ \end{pmatrix}\cdot\begin{pmatrix}Q_{0}\\ \vdots\\ Q_{d}\\ \end{pmatrix}.\;$
@@ 3.3 Trimmed Interpolation
Throughout this section, we show that a polynomial can be (uniquely) interpolated from its evaluations on the trimmed grid points—using essentially the same algorithmic approach as in Section 3.2.
@@
Consider the algorithm TrimmedInterpolation for the proof of Theorem 3.6. As before, we assume that the algorithm has access to the grid $Z$ and for simplicity, we omit $Z$ in the recursive calls.
@@
Given the evaluations $\alpha_{\ell}$ for all $\ell\in\binom{[n]}{\leq D}_{d}$, we show that TrimmedInterpolation correctly interpolates the unique $n$-variate polynomial $P$ by induction on $n$. If $n=0$, simply return the constant $P$ as it does not depend on any variable.
For the remainder of the proof we assume $n\geq 1$. Let $V\coloneqq V(z_{n,0},\dots,z_{n,d})$. It is well-known that $V$ is invertible if and only if the grid points $z_{n,0},\dots,z_{n,d}$ are distinct. Additionally, it implies that the LU decomposition in Line 2, namely $V^{-1}=L\cdot U$, always exists (see e.g. [19]).
In the following, we show that $P(Z_{\ell})=\alpha_{\ell}$ for all $\ell\in\binom{[n]}{\leq D}_{d}$. Note that each $Z_{\ell}$ can be expressed as some $Z_{(\ell^{\prime},j)}$ with $\ell^{\prime}\in\binom{[n-1]}{\leq D-j}_{d}$ and $\sum_{i=1}^{n-1}\ell^{\prime}_{i}+j\leq D$. Fix such a pair $\ell^{\prime}$ and $j$, and let $k=\min\{d,D-\sum_{i=1}^{n-1}\ell^{\prime}_{i}\}$. It holds that
$P\left\lparen Z_{(\ell^{\prime},j)}\right\rparen$ ) $=\sum_{i=0}^{d}z_{n,j}^{i}\cdot P_{i}(Z_{\ell^{\prime}})$
$=\sum_{i=0}^{d}V_{j,i}\cdot\sum_{m=0}^{d}L_{i,m}\cdot Q_{m}(Z_{\ell^{\prime}})$
$=\sum_{i=0}^{d}V_{j,i}\cdot\sum_{m=0}^{d}L_{i,m}\cdot\beta_{(\ell^{\prime},m)}$
$=\sum_{i=0}^{d}V_{j,i}\cdot\sum_{m=0}^{d}L_{i,m}\cdot\sum_{h=0}^{k}U_{m,h}\cdot\alpha_{(\ell^{\prime},h)}$
$=\sum_{i=0}^{d}V_{j,i}\cdot\sum_{h=0}^{k}(V^{-1})_{i,h}\cdot\alpha_{(\ell^{\prime},h)}$
$=\sum_{h=0}^{k}I_{j,h}\cdot\alpha_{(\ell^{\prime},h)}$
$=\alpha_{(\ell^{\prime},j)},$
where $I\in\mathbb{F}^{{d+1}\times{d+1}}$ is the identity matrix. The last equality uses the fact that $j\leq k$. Consequently, TrimmedInterpolation correctly interpolates the unique $n$-variate polynomial $P$ with individual degree $d$ and total degree $D$ such that $P(Z_{\ell})=\alpha_{\ell}$.
As the recursion scheme of TrimmedInterpolation and TrimmedEvaluation are exactly the same, we refer to Lemma 3.4 for a proof of the running time of TrimmedInterpolation. This concludes the proof of Theorem 3.6.

### v2

@@ 3.2 Trimmed Multipoint Evaluation
For the proof of Theorem˜3.1, consider the algorithm TrimmedEvaluation. Throughout we assume that the algorithm has access to the grid $Z$ and for simplicity, we omit $Z$ in the recursive calls.
@@
For $n=0$, the polynomial $P$ does not depend on any variable. Thus, we correctly return the constant $P$. For $n\geq 1$, we write $P(X_{1},\dots,X_{n})=\sum_{i=0}^{d}P_{i}(X_{1},\dots,X_{n-1})\cdot X_{n}^{i}$ to obtain $d+1$ many (n1)(n-1)-variate polynomials $P_{0},\dots,P_{d}$ of degrees $D,\dots,D-d$, respectively. Formally, each polynomial $P_{i}$ is the coefficient of $P$ viewed as a univariate polynomial in $X_{n}$. Let $V\coloneqq V(z_{n,0},\dots,z_{n,d})$ be a Vandermonde matrix. It is well-known that $V$ is invertible if and only if all grid points $z_{n,i}$ are distinct, and in this case the LU decomposition $V=L\cdot U$ in Line 1 always exists (see e.g. [Oruc2000]). Further, let $Q_{0},\dotsc,Q_{d}$ as computed in Line 8 of TrimmedEvaluation. Note that each polynomial $Q_{j}(X_{1},\dotsc,X_{n-1})=\sum_{i=j}^{d}U_{j,i}\cdot P_{i}(X_{1},\dotsc,X_{n-1})$ has degree DjD-j. Indeed, the upper triangular structure of $U$ guarantees that $Q_{j}$ has at most degree DjD-j, because the first non-zero entry in the $j$-th row of $U$ corresponds to the polynomial $P_{j}$ of degree DjD-j while all other $P_{i}$’s with (possible) non-zero coefficients have a smaller degree.
[…]
@@
[…]
@@
Let $T(n,d,D)$ be the running time of TrimmedEvaluation. The time to compute the LU decomposition in Line $4$ and the matrix-vector products in Lines $5$ and $9$ can be bounded by $\operatorname{poly}(n,d)$. Consequently, the running time of the algorithm without the recursive calls in Line $7$ is $\binom{n}{\leq D}_{d}\cdot M(n,d)$ for some function [^6] In our context, we are content with bounding $M(n,d)=\operatorname{poly}(n,d)$; however, let us remark that following van der Hoeven and Schost [HoevenS13] we could achieve a dependence on $d$ that is only polylogarithmic. The main insight is that the L- and U-factors of a Vandermonde matrix are sufficiently structured to support matrix-vector operations in time $O(d(\log d)^{O(1)})$. $M(n,d)=\operatorname{poly}(n,d)$. Thus, the algorithm admits the following recurrence
@@
Combining Lemma˜3.2 and Lemma˜3.4 concludes the proof of Theorem˜3.1.
@@
5 Compute the UL decomposition $V(z_{n,0},\dots,z_{n,d})^{-1}=U\cdot L$.
6 for $\ell^{\prime}\in\binom{[n-1]}{\leq D}_{d}$ do
@@ 7   Let $k=\min\{d,D-\sum_{j=1}^{n-1}\ell^{\prime}_{j}\}$ and compute all evaluations
$\hskip 35.56593pt\begin{pmatrix}\beta_{(\ell^{\prime},0)}\\ \vdots\\ \beta_{(\ell^{\prime},k)}\\ \end{pmatrix}=\begin{pmatrix}L_{0,0}&&\\ \vdots&\ddots&\\ L_{k,0}&\cdots&U_{k,k}\\ \end{pmatrix}\cdot\begin{pmatrix}\alpha_{(\ell^{\prime},0)}\\ \vdots\\ \alpha_{(\ell^{\prime},k)}\\ \end{pmatrix}$.\;
@@
[…]
@@
$\hskip 65.44142pt\begin{pmatrix}P_{0}\\ \vdots\\ P_{d}\\ \end{pmatrix}=\begin{pmatrix}U_{0,0}&\cdots&U_{0,d}\\ &\ddots&\vdots\\ &&U_{d,d}\\ \end{pmatrix}\cdot\begin{pmatrix}Q_{0}\\ \vdots\\ Q_{d}\\ \end{pmatrix}$.\;
@@ 3.3 Trimmed Interpolation
Throughout this section, we show that a polynomial can be (uniquely) interpolated from its evaluations on the trimmed grid points—using essentially the same algorithmic approach as in Section˜3.2.
@@
Consider the algorithm TrimmedInterpolation for the proof of Theorem˜3.6. As before, we assume that the algorithm has access to the grid $Z$ and for simplicity, we omit $Z$ in the recursive calls.
@@
Given the evaluations $\alpha_{\ell}$ for all $\ell\in\binom{[n]}{\leq D}_{d}$, we show that TrimmedInterpolation correctly interpolates the unique $n$-variate polynomial $P$ by induction on $n$. If $n=0$, simply return the constant $P$ as it does not depend on any variable. For the remainder of the proof we assume $n\geq 1$. The induction hypothesis implies that we have $Q_{j}(Z_{\ell}^{\prime})=\beta_{(\ell^{\prime},j)}$ as defined in Line 2, for all $\ell^{\prime}\in\binom{[n-1]}{\leq D-j}_{d}$.
Let $V\coloneqq V(z_{n,0},\dots,z_{n,d})$. It is well-known that $V$ is invertible if and only if the grid points $z_{n,0},\dots,z_{n,d}$ are distinct. Moreover, in this case the UL decomposition $V^{-1}=U\cdot L$ from Line 2 always exists (see e.g. [Oruc2000]). As the matrices $U$ and $L$ are invertible, too, in the following calculation we will use the equivalent formulation $V\cdot U=L^{-1}$.
We now show that $P(Z_{\ell})=\alpha_{\ell}$ for all $\ell\in\binom{[n]}{\leq D}_{d}$. Note that each $Z_{\ell}$ can be expressed as some $Z_{(\ell^{\prime},j)}$ with $\ell^{\prime}\in\binom{[n-1]}{\leq D-j}_{d}$ and $\sum_{i=1}^{n-1}\ell^{\prime}_{i}+j\leq D$. Fix such a pair $\ell^{\prime}$ and $j$. As in the algorithm, let $k=\min\{d,D-\sum_{i=0}^{n-1}\ell_{i}^{\prime}\}$. Clearly we have $j\leq k$. It holds that:
[…]
$=\sum_{i=0}^{d}V_{j,i}\cdot\sum_{m=0}^{d}U_{i,m}\cdot Q_{m}(Z_{\ell^{\prime}})$
$=\sum_{i=0}^{d}V_{j,i}\cdot\sum_{m=i}^{d}U_{i,m}\cdot Q_{m}(Z_{\ell^{\prime}})$
$=\sum_{m=0}^{d}Q_{m}(Z_{\ell^{\prime}})\cdot\sum_{i=0}^{m}V_{j,i}\cdot U_{i,m}$
$=\sum_{m=0}^{d}Q_{m}(Z_{\ell^{\prime}})\cdot(L^{-1})_{j,m}$
$=\sum_{m=0}^{j}Q_{m}(Z_{\ell^{\prime}})\cdot(L^{-1})_{j,m}$
$=\sum_{m=0}^{j}\beta_{(\ell^{\prime},m)}\cdot(L^{-1})_{j,m}$
$=\sum_{m=0}^{k}\beta_{(\ell^{\prime},m)}\cdot(L^{-1})_{j,m}$
$=\alpha_{(\ell^{\prime},j)}.$
Consequently, TrimmedInterpolation correctly interpolates the unique $n$-variate polynomial $P$ with individual degree $d$ and total degree $D$ such that $P(Z_{\ell})=\alpha_{\ell}$.
As the recursion scheme of TrimmedInterpolation and TrimmedEvaluation are exactly the same, we refer to Lemma˜3.4 for a proof of the running time of TrimmedInterpolation. This concludes the proof of Theorem˜3.6.